OpenStudy (lgbasallote):

a tank contains 100 gal of water. brine enters the tank at the rate of 3 gpm. the mixture, thoroughly stirred, leaves the tank at the rate of 2 gpm. if the concentration of the brine at the end of 20 minutes is to be 2 lb/gal, what should be the concentration of the brine entering the tank?

OpenStudy (lgbasallote):

@amistre64 would you happen to know this cheese?

OpenStudy (amistre64):

i think i recall it

OpenStudy (lgbasallote):

yay!

OpenStudy (amistre64):

in-out = total amount that stays right?

OpenStudy (lgbasallote):

i believe so?

OpenStudy (zepp):

42.

OpenStudy (lgbasallote):

input - output = accumulation

OpenStudy (lgbasallote):

does that mean same thing?

OpenStudy (amistre64):

yes

OpenStudy (lgbasallote):

i see i see

OpenStudy (amistre64):

how do we calculate salinity?

OpenStudy (amistre64):

is the brine 100% coming in?

OpenStudy (lgbasallote):

what does salinity mean?

OpenStudy (amistre64):

brine is salt water; salt is a saline solution ....

OpenStudy (lgbasallote):

well the amt of brine coming in is represented by c1 the amt of water coming in is r1

OpenStudy (lgbasallote):

so c1 is 100% salt i believe

OpenStudy (amistre64):

you have lbs/gal to figure out, but i dont see where you can determine how many lbs there is coming and going to start of with

OpenStudy (anonymous):

See this http://tutorial.math.lamar.edu/Classes/DE/Modeling.aspx

OpenStudy (lgbasallote):

here is the formula \[\large r_1 c_1 - r_2 [\frac{Q}{V_o + (r_1 - r_2)t}] = \frac{dQ}{dt}\]

OpenStudy (amistre64):

after 20 minutes tho, i would assume there is about 120 gallons in the tank; but i cant see a way of calculating lbs to gallons without knowing the ratio of the incoming brine

OpenStudy (lgbasallote):

i remember my teacher doing something like linear DE...does that help?

OpenStudy (amistre64):

it might, but doesnt jog any memories :) at any rate, i have to get back to pulling files for statistical analysises. Good luck :)

OpenStudy (lgbasallote):

uhhmm okay..good luck to you too :)

OpenStudy (lgbasallote):

well i guess no one can solve this here.....